OpenStudy (anonymous):

What is the real value of "sin"? I mean, when I hit the "sin" button on the calculator and then hit the "equal" button it shows an amount like, -1.1270x10^09, but then I repeat it and it shows another value. :S

OpenStudy (anonymous):

sin doesn't have a value on its own. When you try to calculate \(sin\theta\), what the calculator does is find the length of the 'opposite' side of a right-angled triangle with the angle \(\theta\), when the hypotenuse is length 1. When you press 'sin' and 'equals', your calculator probably just takes the last calculated value, and makes it the angle. For example, you might have typed in \(6 \times 3 =\). When you press 'sin' 'equals', the calculator will calculate \(sin(18)\).

OpenStudy (anonymous):

In the exams that we're about to take, calculators are not allowed, so to get the value of sin, we multiply it by one?

OpenStudy (anonymous):

Do you have access to log tables, or maths tables?

OpenStudy (anonymous):

Nope and nope.

OpenStudy (anonymous):

I have no idea how you're supposed to calculate it then I'm afraid. We've always been allowed to at least use maths tables... It's possible that in the exam, you'll only be given certain angles. sin takes on nice values in certain places. \(\theta\) 30 45 60 90 180 sin\(\theta\) \(\frac{1}{2}\) \(\frac{1}{\sqrt{2}}\) \(\frac{\sqrt{3}}{2}\) 1 0

OpenStudy (anonymous):

Will keep that in mind. Atleast now I have some values without using the calculator. Thanks! :)

OpenStudy (anonymous):

There are some similar values for cos and tan. Look them up somewhere!

OpenStudy (anonymous):

Lol thanks. :)

OpenStudy (anonymous):

i think i read somewhere that calculators work out trig ratios using mclaurins series

OpenStudy (anonymous):

U just have to remember a couple basic triangles (eg 1,1,sqrt 2)and Pythagorus, u can work them out